OpenStudy (anonymous):

ok so write the slope intercent form of the equation of the line through the given points.through:(3,3) and (0,2) i think the slope is -1/3

OpenStudy (anonymous):

slope would be 1/3 either 3 - 2 over 3 - 0 or 2 - 3 over 0 - 3 both are 1/3

OpenStudy (anonymous):

you can use y2-y1/x2-x1 or vice versa as long as ys are on top

OpenStudy (anonymous):

use m=1/3 and either point and plug into to y=mx+b and solve for b then plug m and b into equation

OpenStudy (anonymous):

k so i think y=1/3x+2

OpenStudy (anonymous):

GREAT JOB!

OpenStudy (anonymous):

:) yay thank u!

OpenStudy (anonymous):

welcome...you can always check by plugging both points into your equation to see if its it true...if the solutions are true that means the lines passes through those points